OpenStudy (dls):

A bomb of mass 10kg is projected from a point time t=0,theta=37 degrees, above the horizontal with a speed =30m/s.At some point after the projection it explodes into 2 fragments,one fragment of mass 4kg is observed at (105,43)m at time t=2s.Then the position of 2nd fragment at t=2s is?

OpenStudy (dls):

@experimentX

OpenStudy (experimentx):

give it a try ... find the position of bomb without exploding at t=2s

OpenStudy (dls):

I have to use this right? \[\frac{m _{1}x _{1}+m _{2}x _{2}}{m1+m2}\]

OpenStudy (dls):

for the X position

OpenStudy (experimentx):

yeah ... center of mass will be at the position of the bomb without exploding. first find the CM.

OpenStudy (dls):

Initially it will be 0?

OpenStudy (experimentx):

no ... find the cm at 2s .. not 0s

OpenStudy (dls):

\[\frac{4 \times 105 + 6 \times x2}{10} \]

OpenStudy (dls):

for X?

OpenStudy (dls):

or do i have to equate to 105 itslf?

OpenStudy (dls):

4x1+6x2/10=105?

OpenStudy (experimentx):

this is just projectile motion ... what would be the position of the stone (not bomb) after 2s? with same conditions ??

OpenStudy (dls):

90 sin74?

OpenStudy (dls):

@experimentX ?

OpenStudy (experimentx):



OpenStudy (dls):

what i found is range